# Simultaneous FUV and NUV observations of T Tauri stars with UVIT/AstroSat: probing accretion process in young stars

### v1

@@ Abstract
We present results from simultaneous FUV and NUV observations of T-Tauri stars (TTSs) in the Taurus molecular cloud with UVIT/AstroSat. This is the very first UVIT study of TTSs. From the spectral energy distribution of TTSs from FUV to IR, we show that classical TTSs (CTTSs) emit significantly higher UV excess compared to weak-line TTSs (WTTSs). The equivalent black-body temperatures corresponding to the UV excess in CTTSs ( K) are also found to be relatively higher than that in WTTSs ( K). From the UV excess, we have re-classified two WTTSs (BS Tau, V836 Tau) as CTTSs, which has been supported by the follow-up optical spectroscopic study using the Himalayan Chandra Telescope (HCT), showing strong H line emission. We find that CTTS show strong excess emission in both FUV (10) and NUV (10) bands, while WTTS show strong excess only in the FUV (10), suggesting that excess emission in NUV can be used as a tool to classify the TTSs. We also find a linear correlation between UV luminosity (a primary indicator of mass accretion) and H luminosity (a secondary indicator of mass accretion) with a slope of 1.210.17 and intercept of 2.180.56.
@@ 1 Introduction
T Tauri stars (TTS) are low-mass pre-main-sequence (PMS) stars, generally categorized into Classical TTS (CTTS) and Weak-line TTS (WTTS) based on their strength of H emission (Alcala et al., 1993; Duvert et al., 2000; Gras-Velázquez & Ray, 2005). CTTS show strong and broad H emission indicating active ongoing accretion from the circumstellar disk onto the central star, while WTTS show weak and narrow H emission, suggesting weak or no accretion (Herbig & Bell, 1988; Martín, 1998; White & Basri, 2003; Barrado y Navascués & Martín, 2003).
@@
Circumstellar disks form during the process of star formation due to the conservation of angular momentum and in their short lifetimes ( a few Myrs) aid both star and planet formation. Material from the disk around the CTTS is channelled onto the stars along the strong magnetic field lines and generates accretion shocks on the stellar surface, producing hot spots (Hartmann et al., 2016, and references therein). The kinetic energy of the freely in-falling material is dissipated in these accretion shocks. Behind these shocks, the energy is converted into radiation which flows back out through the shocks and into the in-falling material. Most of the escaping radiation peaks in the ultraviolet with estimated black body equivalent temperatures of 10 K (Calvet & Gullbring, 1998). Thus, accreting stars are characterized by strong UV excesses (Gorti & Hollenbach, 2009). The cool and dusty layers of the disk absorb this radiation and re-emit at the infrared wavelengths. Hence, the accretion process causes the release of excess energy not only in the UV but also in the IR regions of the spectral energy distribution of a CTTS. The UV emissions mainly originate from the hot spots produced by accretion shocks, while near-IR (NIR) and mid-IR emissions come from smaller disk radii while far-IR comes from the midplane of the outer disk. As the disk disperses, the SED gradually shows less excess in the IR region (Lada, 1987). Non-accreting WTTS also shows excess in UV due to chromospheric activity and significantly reduced excess in IR due to the presence of a smaller or depleted disk (Ingleby et al., 2011; Schneider et al., 2020).
@@
FUV photons from the star play an important role in heating disk gas and can drive massive thermal winds that eventually deplete disk material. Recent research suggests that FUV-aided dispersal and heating can have consequences for angular momentum transport through magneto-centrifugal winds (e.g., Bai et al. (2016)) and the formation of planetesimals (Carrera et al., 2017). FUV photons significantly affect disk chemistry—the differences in the molecule content of low-mass TTS and intermediate-mass Herbig stars is believed to be due to the stronger UV flux from the latter (Pascucci et al., 2014). This could have implications for the composition of planetesimals and planets that eventually form in these disks. Even after the disk is dispersed, the bright FUV flux from WTTS could impact the evolution of proto-planets and potentially strip their atmospheres (Owen & Wu, 2016). Characterizing FUV emission from T Tauri stars and understanding the time evolution of the spectrum is critical to understanding disk evolution (Gorti & Hollenbach, 2009), and more fundamentally, the accretion process itself.
@@
There have been several UV spectral surveys of TTS by the IUE (Valenti et al., 2003) and the STIS instruments on the HST (Yang et al., 2012). While these studies have yielded valuable information on the FUV spectrum, most of the measurements of the FUV and NUV excesses are not simultaneous. It is thought that the NUV and the FUV flux originate from different regions of the accretion flow. The NUV Balmer continuum arises from the dense pre-shock region of the accretion column and the FUV most likely arises near the hot spot (Ulrich, 1976; Calvet & Gullbring, 1998; Ingleby et al., 2013; Hartmann et al., 2016); this makes contemporaneous observations of both bands necessary for understanding the link between the NUV and the FUV emitting regions and the nature of the accretion process (France et al., 2014). UVIT on AstroSat, with its capability of accurate and simultaneous multi-band UV photometry, is the ideal instrument for probing the FUV-emitting regions around young stars. The use of multiple filters, three filters each in the FUV ( 1481 Å, 1541 Å & 1608 Å) and NUV ( 2447 Å, 2632 Å & 2792 Å) bands, will give us three-wavelength points allowing us to better reconstruct the flux distribution across FUV and NUV regions.
@@
In this paper, we present simultaneous multi-band observations in FUV and NUV bands of T-Tauri Stars in the Taurus Molecular Cloud. We have modelled the UV flux with a black-body spectrum and have determined the corresponding temperature and luminosity. For CTTS, the black-body luminosity will provide a direct measure of accretion luminosity. We investigated if there is any significant difference in the black-body temperatures between CTTS and WTTS, i.e., between peak temperature for the emission due to accretion and emission due to chromospheric activity. We have also revisited the classification scheme based on the amount of UV excess emitted by TTS over photospheric emission and the strength of the H emission. We also search for correlations between UV luminosity and other commonly used accretion/disk tracers such as H luminosity and IR excess emission.
@@
The remaining sections of this paper are arranged as follows. In Sect. 2, we discuss the UVIT data, the follow-up optical spectroscopic observation from ground-based telescopes, and SED analysis of T-Tauri stars. In Sect. 3, we present our results and discuss them. In the Sect. 4, we summarize our results from this study and their implications.
@@ 2.1 Photometry: UVIT observations
Four TTS fields in the Taurus Molecular Cloud (TMC) were observed with UVIT onboard AstroSat (proposal id: A04-210; PI: Annapurni Subramaniam). Each of these fields has a diameter of 28′ and were centered on the following TTS: FM Tau, V836 Tau, BS Tau, and HD283782. All the observations were carried out using three FUV and three NUV filters in January of 2018. The names of the FUV and NUV filters and corresponding exposure times for all targets are listed in Table 1. FM Tau is observed in the narrow band filter, N279N, centered at the Mg-II line, while other TTSs are observed in the N242W wide band, keeping the other two medium bands (N245M and N263M) as common filters for all the targets.
The observations were completed in multiple orbits. We applied corrections for spacecraft drift, flat-field, and distortion using the software CCDLAB (Postma & Leahy, 2017) and created images for each orbit. Then, the orbit-wise images were co-aligned and combined to generate final science-ready images in each filter. Astrometry was also performed using CCDLAB by comparing the Gaia-DR2 source catalogue. The science-ready images were created for an area of 4K4K in size with a scale of 0.416/pixel. The details about the telescope and the instruments are available in Subramaniam et al. (2016); Tandon et al. (2017a) and the instrument calibration can be found in Tandon et al. (2017b, 2020).
@@
Given the large field of view (FOV) of UVIT, it is possible that other TTS can also be present in the same FOV. To detect these serendipitous TTS, we cataloged UVIT-detected stars in the various fields and cross-matched them with the Gaia and UV catalogs of the TMC members from Esplin & Luhman (2019), Nayak et al. (2023) and Inés Gómez de Castro et al. (2023) to search for more UV counterparts to TMC members. With UVIT we were able to detect seven more TTS in the field of FM Tau. These are V773 Tau, CW Tau, FN Tau, FO Tau, CIDA 1, Anon 1, and 2MASS J04141188+2811535. While three of the targets (V773 Tau, CW Tau, FN Tau) were detected in both NUV and FUV, four of the TTSs (FO Tau, CIDA 1, Anon 1, and 2MASS J04141188+2811535; hereafter J0414) were only detected in NUV. No additional TTSs were detected in the FOV of BS Tau, V836 Tau and HD 283782. The UVIT magnitudes in various filters of all the 11 detected TTSs and their coordinates are given in Table 1.
@@ 2.2 Spectroscopy: HCT observation
We obtained low-resolution optical spectroscopic of all 11 TTSs. As the optical spectroscopy and UV observations are not simultaneous, it is necessary to obtain multi-epoch spectra of these sources, which will allow us to identify if there is any variability in the H emission, i.e. variability in the accretion rate. We observed five of our targets with Himalayan Faint Object Spectrograph Camera (HFOSC)222Further details of the instruments and telescopes are available at http://www.iiap.res.in/iao/hfosc.html mounted on the 2 m Himalayan Chandra Telescope (HCT). We also searched the LAMOST database and found optical spectra of eight out of eleven of our targets. In Table 2, we have listed the names of the sources, from where the spectra are obtained, and observation dates of different epochs for each source. Except for Anon1 and HD283782, we obtain multi-epoch observations of all the other sources.
For the HCT spectra, the wavelength range was covered using Grism 8 (5500–9000Å) with an effective resolving power of 1050. Flat frames have been taken before each on-target observation, while bias frames have been taken both before and after the on-target observation. The FeNe lamp spectra were also obtained after each on-target observation for wavelength calibration. The HCT spectra were reduced in a standard procedure after bias subtraction and flat-field correction using the standard tasks in the IRAF software and the HCT pipeline HAPLI (Narang et al., 2023). Then wavelength calibration is performed on extracted spectra for further analysis.
@@
| N245M | 1221 | 17.46 | F148W | 1208 | 19.98 | ||||
| FM Tau | 4 14 13.6 | 28 12 49.2 | N263M | 958 | 16.98 | F154W | 1207 | 19.75 | CTTS |
| N279N | 1211 | 16.56 | F169M | 945 | 19.62 | ||||
| N245M | 1221 | 17.80 | F148W | 1208 | 20.65 | ||||
| V773 Tau* | 4 14 13.6 | 28 12 49.2 | N263M | 958 | 16.92 | F154W | 1207 | 19.94 | WTTS |
| N279N | 1211 | 15.95 | F169M | 945 | 20.0 | ||||
| N245M | 1221 | 16.83 | F148W | 1208 | 19.38 | ||||
| CW Tau* | 04 14 17.0 | 28 10 57.8 | N263M | 958 | 16.24 | F154W | 1207 | 19.19 | CTTS |
| N279N | 1211 | 15.32 | F169M | 945 | 18.88 | ||||
| N245M | 1221 | 19.77 | F148W | 1208 | |||||
| FO Tau* | 04 14 49.3 | 28 12 30.46 | N263M | 958 | 19.05 | F154W | 1207 | CTTS | |
| N279N | 1211 | 17.95 | F169M | 945 | |||||
| N245M | 1221 | 18.64 | F148W | 1208 | 21.02 | ||||
| FN Tau* | 04 14 14.6 | 28 27 58.1 | N263M | 958 | F154W | 1207 | 20.85 | CTTS | |
| N279N | 1211 | 16.72 | F169M | 945 | 20.39 | ||||
| N245M | 1221 | 21.58 | F148W | 1188 | |||||
| CIDA 1* | 04 14 17.520 | 28 06 9.0 | N263M | 958 | 20.61 | F154W | 1207 | CTTS | |
| N279N | 1211 | 20.08 | F169M | 945 | |||||
| N245M | 1221 | 21.48 | F148W | 1188 | |||||
| J041411.88 | 04 14 11.88 | 28 11 53.31 | N263M | 958 | 20.60 | F154W | 1207 | CTTS | |
| +281153.3* | N279N | 1211 | 19.63 | F169M | 945 | ||||
| N245M | 1221 | 21.92 | F148W | 1188 | |||||
| Anon 1* | 04 13 27.216 | 28 16 22.8 | N263M | 958 | 20.82 | F154W | 1207 | WTTS | |
| N279N | 1211 | 19.63 | F169M | 945 | |||||
| N242W | 1214 | 17.92 | F148W | 1188 | 19.44 | ||||
| BS Tau | 4 58 51.4 | 28 31 24.2 | N245M | 1228 | 18.11 | F154W | 1190 | 19.24 | CTTS |
| N263M | 1225 | 17.83 | F169M | 1189 | 19.25 | ||||
| N242W | 1228 | 19.54 | F148W | 1199 | 20.68 | ||||
| V836 Tau | 5 3 6.6 | 25 23 19.7 | N245M | 1227 | 19.68 | F154W | 1200 | 20.32 | CTTS |
| N263M | 1205 | 19.05 | F169M | 1187 | 20.18 | ||||
| N242W | 1209 | 16.20 | F148W | 1193 | 19.08 | ||||
| HD283782 | 4 44 54.4 | 27 17 45.2 | N245M | 1228 | 15.98 | F154W | 1207 | 18.94 | WTTS |
| N263M | 860 | 15.05 | F169M | 824 | 18.91 |
@@
Note. — *In the UVIT field of FM Tau, we found seven more TTS candidates having similar exposure times.
| Name | instruments | observation date | EW(H) | mean EW(H) | L | Classification | |
| () | (10) | ||||||
| Anon 1 | LAMOST | 2014-11-05 | -3.9 | 3.9 | — | — | WTTS |
| BS Tau | HCT | 2021-01-17 | -31.1 | 33.12.0 | 0.0110.001 | 2.17 | CTTS |
| 2021-09-06 | -35.1 | ||||||
| CIDA 1 | LAMOST | 2012-01-04 | -141.9 | 205.5 | 0.004 | 0.70 | CTTS |
| 2011-12-18 | -224.8 | ||||||
| 2012-01-13 | -133.3 | ||||||
| 2012-01-22 | -97.7 | ||||||
| 2013-02-08 | -271.7 | ||||||
| 2014-01-25 | -242.1 | ||||||
| 2015-10-13 | -231.5 | ||||||
| 2015-12-28 | -273.2 | ||||||
| 2016-12-31 | -232.9 | ||||||
| CW Tau | LAMOST | 2014-01-25 | -87.1 | 126.8 | 0.54 | 107.69 | CTTS |
| 2014-01-30 | -232.2 | ||||||
| 2014-11-05 | -60.4 | ||||||
| FM Tau | LAMOST | 2014-01-30 | -112.8 | 118.8 | 0.083 | 16.75 | |
| HCT | 2021-01-17 | -100.7 | |||||
| 2021-09-06 | -142.9 | ||||||
| FN Tau | LAMOST | 2014-01-25 | -26.9 | 26.0 | 0.022 | 4.32 | CTTS |
| 2014-01-30 | -21.8 | ||||||
| 2014-11-05 | -29.4 | ||||||
| FO Tau | LAMOST | 2011-12-18 | -136.4 | 143.5 | 0.059 | 11.68 | CTTS |
| 2012-01-04 | -172.7 | ||||||
| 2012-01-13 | -135.9 | ||||||
| 2013-02-08 | -135.2 | ||||||
| 2014-01-25 | -137.4 | ||||||
| 2014-01-30 | -158.2 | ||||||
| HD283782 | HCT | 2021-09-06 | -1.9 | 1.9 | — | — | WTTS |
| V773 Tau | LAMOST | 2014-11-05 | -3.2 | 2.5 | — | — | WTTS |
| HCT | 2021-09-06 | -1.9 | |||||
| V836 Tau | HCT | 2021-01-17 | -8.9 | 13.3 | 0.013 | 2.60 | CTTS |
| 2021-09-06 | -17.6 | ||||||
| J041411.88 | LAMOST | 2012-01-22 | -319.4 | 357.4 | 0.002 | 0.49 | CTTS |
| 2012-01-23 | -404.6 | ||||||
| 2015-12-28 | -348.2 |
Note. —

### v2

@@ Abstract
We present results from simultaneous FUV and NUV observations of T-Tauri stars (TTSs) in the Taurus molecular cloud with UVIT/AstroSat. This is the very first UVIT study of TTSs. From the spectral energy distribution of TTSs from FUV to near-IR, we show that classical TTSs (CTTSs) emit significantly higher UV excess compared to weak-line TTSs (WTTSs). The equivalent black-body temperatures corresponding to the UV excess in CTTSs ( K) are also found to be relatively higher than that in WTTSs ( K). From the UV excess, we have re-classified two WTTSs (BS Tau, V836 Tau) as CTTSs, which has been supported by the follow-up optical spectroscopic study using the Himalayan Chandra Telescope (HCT), showing strong H line emission. We find that CTTSs show strong excess emission in both FUV (107) and NUV (103) bands, while WTTSs show strong excess only in the FUV (105), suggesting that excess emission in NUV can be used as a tool to classify the TTSs. We also find a linear correlation between UV luminosity (a primary indicator of mass accretion) and H luminosity (a secondary indicator of mass accretion) with a slope of 1.200.22 and intercept of 2.160.70.
@@ 1 Introduction
T Tauri stars (TTSs) are low-mass pre-main-sequence (PMS) stars, generally categorized into Classical TTSs (CTTSs) and Weak-line TTSs (WTTSs) based on their strength of H emission (Alcala et al., 1993; Duvert et al., 2000; Gras-Velázquez & Ray, 2005). CTTSs show strong and broad H emission indicating active ongoing accretion from the circumstellar disk onto the central star, while WTTSs show weak and narrow H emission, suggesting weak or no accretion (Herbig & Bell, 1988; Martín, 1998; White & Basri, 2003; Barrado y Navascués & Martín, 2003).
@@
Circumstellar disks form during the process of star formation due to the conservation of angular momentum and in their short lifetimes ( a few Myrs) aid both star and planet formation. Material from the disk around the CTTSs is channelled onto the stars along the strong magnetic field lines and generates accretion shocks on the stellar surface, producing hot spots (Hartmann et al., 2016, and references therein). The kinetic energy of the freely in-falling material is dissipated in these accretion shocks. Behind these shocks, the energy is converted into radiation which flows back out through the shocks and into the in-falling material. Most of the escaping radiation peaks in the ultraviolet with estimated black body equivalent temperatures of 104 K (Calvet & Gullbring, 1998). Thus, accreting stars are characterized by strong UV excesses (Gorti & Hollenbach, 2009). The cool and dusty layers of the disk absorb this radiation and re-emit at the infrared wavelengths. Hence, the accretion process causes the release of excess energy not only in the UV but also in the IR regions of the spectral energy distribution of a CTTS. The UV emissions mainly originate from the hot spots produced by accretion shocks, while near-IR (NIR) and mid-IR emissions come from smaller disk radii while far-IR comes from the midplane of the outer disk. As the disk disperses, the SED gradually shows less excess in the IR region (Lada, 1987). Non-accreting WTTS also shows excess in UV due to chromospheric activity and significantly reduced excess in IR due to the presence of a smaller or depleted disk (Ingleby et al., 2011; Schneider et al., 2020).
@@
FUV photons from the star play an important role in heating disk gas and can drive massive thermal winds that eventually deplete disk material. Recent research suggests that FUV-aided dispersal and heating can have consequences for angular momentum transport through magneto-centrifugal winds (e.g., Bai et al. (2016)) and the formation of planetesimals (Carrera et al., 2017). FUV photons significantly affect disk chemistry—the differences in the molecule content of low-mass TTSs and intermediate-mass Herbig stars are believed to be due to the stronger UV flux from the latter (Pascucci et al., 2014). This could have implications for the composition of planetesimals and planets that eventually form in these disks. Even after the disk is dispersed, the bright FUV flux from WTTS could impact the evolution of (proto-) planets and potentially strip their atmospheres (Owen & Wu, 2016). Characterizing FUV emission from TTSs and understanding the time evolution of the spectrum is critical to understanding disk evolution (Gorti & Hollenbach, 2009), and more fundamentally, the accretion process itself.
@@
There have been several UV spectral surveys of TTSs by the IUE (Valenti et al., 2003) and the STIS instruments on the HST (Yang et al., 2012). While these studies have yielded valuable information on the FUV spectrum, most of the measurements of the FUV and NUV excesses are not simultaneous. It is thought that the NUV and the FUV flux originate from different regions of the accretion flow. The NUV Balmer continuum arises from the dense pre-shock region of the accretion column and the FUV most likely arises near the hot spot (Ulrich, 1976; Calvet & Gullbring, 1998; Ingleby et al., 2013; Hartmann et al., 2016); this makes contemporaneous observations of both bands necessary for understanding the link between the NUV and the FUV emitting regions and the nature of the accretion process (France et al., 2014; Nayak et al., 2024). UVIT on AstroSat, with its capability of accurate and simultaneous multi-band UV photometry, is the ideal instrument for probing the FUV-emitting regions around young stars. The use of multiple filters, three filters each in the FUV ( 1481 Å, 1541 Å & 1608 Å) and NUV ( 2447 Å, 2632 Å & 2792 Å) bands, will give us three-wavelength points allowing us to better reconstruct the flux distribution across FUV and NUV regions.
@@
In this paper, we present simultaneous multi-band observations in FUV and NUV bands of T-Tauri Stars in the Taurus Molecular Cloud. We have modelled the UV flux with a black-body spectrum and have determined the corresponding temperature and luminosity. For a CTTS, the black-body luminosity will provide a direct measure of accretion luminosity. We investigated if there is any significant difference in the black-body temperatures between CTTSs and WTTSs, i.e., between peak temperature for the emission due to accretion and emission due to chromospheric activity. We have also revisited the classification scheme based on the amount of UV excess emitted by TTSs over photospheric emission and the strength of the H emission. We also search for correlations between UV luminosity and other commonly used accretion/disk tracers such as H luminosity and NIR excess emission.
@@
The remaining sections of this paper are arranged as follows. In Sect. 2, we discuss the UVIT data, the follow-up optical spectroscopic observation from ground-based telescopes, and the SED analysis of T-Tauri stars. In Sect. 3, we present our results and discuss them. In the Sect. 4, we summarize our results from this study and their implications.
@@ 2.1 Photometry: UVIT observations
Four TTSs fields in the Taurus Molecular Cloud (TMC) were observed with UVIT onboard AstroSat (proposal id: A04-210; PI: Annapurni Subramaniam). Each of these fields has a diameter of 28′ and was centered on the following TTS: FM Tau, V836 Tau, BS Tau, and HD283782. All the observations were carried out using three FUV and three NUV filters in January 2018. The names of the FUV and NUV filters and corresponding exposure times for all targets are listed in Table 1. FM Tau is observed in the narrow band filter, N279N, centered at the Mg-II line, while other TTSs are observed in the N242W wide band, keeping the other two medium bands (N245M and N263M) as common filters for all the targets.
The observations were completed in multiple orbits. We applied corrections for spacecraft drift, flat-field, and distortion using the software CCDLAB (Postma & Leahy, 2017) and created images for each orbit. Then, the orbit-wise images were co-aligned and combined to generate final science-ready images in each filter. Astrometry was also performed using CCDLAB by comparing the Gaia-DR2 source catalog. The science-ready images were created for an area of 4K4K in size with a scale of 0.′′416/pixel. The details about the telescope and the instruments are available in Subramaniam et al. (2016); Tandon et al. (2017a) and the instrument calibration can be found in Tandon et al. (2017b, 2020).
@@
Given the large field of view (FOV) of UVIT, it is possible that other TTSs can also be present in the same FOV. To detect these serendipitous TTSs, we cataloged UVIT-detected stars in the various fields and cross-matched them with the Gaia and UV catalogs of the TMC members from Esplin & Luhman (2019), Nayak et al. (2023) and Inés Gómez de Castro et al. (2023) to search for more UV counterparts to TMC members. With UVIT we were able to detect six more TTSs in the field of FM Tau. These are V773 Tau, CW Tau, FO Tau, CIDA 1, Anon 1, and 2MASS J04141188+2811535. While two of the targets (V773 Tau and CW Tau) were detected in both NUV and FUV, four of the TTSs (FO Tau, CIDA 1, Anon 1, and 2MASS J04141188+2811535; hereafter J0414) were only detected in NUV. No additional TTSs were detected in the FOV of BS Tau, V836 Tau and HD 283782. The final UVIT magnitudes along with the corresponding photometric errors in various filters of all the 10 detected TTSs and their coordinates are listed in Table 1.
@@ 2.2 Spectroscopy: HCT observation
We obtained low-resolution optical spectroscopic of all 10 TTSs. As the optical spectroscopy and UV observations are not simultaneous, it is necessary to obtain multi-epoch spectra of these sources, which will allow us to identify if there is any variability in the H emission, i.e. variability in the accretion rate. We observed five of our targets with Himalayan Faint Object Spectrograph Camera (HFOSC)222Further details of the instruments and telescopes are available at http://www.iiap.res.in/iao/hfosc.html mounted on the 2 m Himalayan Chandra Telescope (HCT). We also searched the LAMOST database (Cui et al., 2012) and found optical spectra of eight out of eleven of our targets. In Table 2, we have listed the names of the sources, from where the spectra are obtained, and observation dates of different epochs for each source. We obtained multi-epoch observations for all of the sources except Anon1 and HD283782.
For the HCT spectra, the wavelength range was covered using Grism 8 (5500–9000Å) with an effective resolving power of 1050. Flat frames were taken before each on-target observation, while bias frames were taken both before and after the on-target observation. The FeNe lamp spectra were also obtained after each on-target observation for wavelength calibration. The HCT spectra were reduced in a standard procedure after bias subtraction and flat-field correction using the standard tasks in the IRAF software and the HCT pipeline HAPLI (Narang et al., 2023). Then wavelength calibration was performed on extracted spectra for further analysis.
@@
| N245M | 1221 | 17.460.02 | F148W | 1208 | 19.980.08 | ||||
| FM Tau | 4 14 13.6 | 28 12 49.2 | N263M | 958 | 16.980.02 | F154W | 1207 | 19.750.09 | CTTS |
| N279N | 1211 | 16.560.04 | F169M | 945 | 19.620.11 | ||||
| N245M | 1221 | 17.800.03 | F148W | 1208 | 20.650.11 | ||||
| V773 Tau* | 4 14 13.6 | 28 12 49.2 | N263M | 958 | 16.920.02 | F154W | 1207 | 19.940.10 | WTTS |
| N279N | 1211 | 15.950.03 | F169M | 945 | 20.00.13 | ||||
| N245M | 1221 | 16.830.02 | F148W | 1208 | 19.380.06 | ||||
| CW Tau* | 04 14 17.0 | 28 10 57.8 | N263M | 958 | 16.240.02 | F154W | 1207 | 19.190.07 | CTTS |
| N279N | 1211 | 15.320.02 | F169M | 945 | 18.880.08 | ||||
| N245M | 1221 | 19.770.07 | F148W | 1208 | |||||
| FO Tau* | 04 14 49.3 | 28 12 30.46 | N263M | 958 | 19.050.05 | F154W | 1207 | CTTS | |
| N279N | 1211 | 17.950.07 | F169M | 945 | |||||
| N245M | 1221 | 21.580.15 | F148W | 1188 | |||||
| CIDA 1* | 04 14 17.520 | 28 06 9.0 | N263M | 958 | 20.610.13 | F154W | 1207 | CTTS | |
| N279N | 1211 | 20.080.17 | F169M | 945 | |||||
| N245M | 1221 | 21.480.15 | F148W | 1188 | |||||
| J041411.88 | 04 14 11.88 | 28 11 53.31 | N263M | 958 | 20.600.13 | F154W | 1207 | CTTS | |
| +281153.3* | N279N | 1211 | 19.630.14 | F169M | 945 | ||||
| N245M | 1221 | 21.920.18 | F148W | 1188 | |||||
| Anon 1* | 04 13 27.216 | 28 16 22.8 | N263M | 958 | 20.820.14 | F154W | 1207 | WTTS | |
| N279N | 1211 | 19.630.13 | F169M | 945 | |||||
| N242W | 1214 | 17.920.01 | F148W | 1188 | 19.440.06 | ||||
| BS Tau | 4 58 51.4 | 28 31 24.2 | N245M | 1228 | 18.110.03 | F154W | 1190 | 19.240.06 | CTTS |
| N263M | 1225 | 17.830.03 | F169M | 1189 | 19.250.07 | ||||
| N242W | 1228 | 19.540.03 | F148W | 1199 | 20.680.1 | ||||
| V836 Tau | 5 3 6.6 | 25 23 19.7 | N245M | 1227 | 19.680.06 | F154W | 1200 | 20.320.1 | CTTS |
| N263M | 1205 | 19.050.05 | F169M | 1187 | 20.180.11 | ||||
| N242W | 1209 | 16.200.01 | F148W | 1193 | 19.080.05 | ||||
| HD283782 | 4 44 54.4 | 27 17 45.2 | N245M | 1228 | 15.980.01 | F154W | 1207 | 18.940.05 | WTTS |
| N263M | 860 | 15.050.01 | F169M | 824 | 18.910.08 |
@@
Note. — *In the UVIT field of FM Tau, we found six more TTS candidates having similar exposure times.
| Source | instruments | observation date | EW(H) | Lacc | Classification | ||
| () | L⊙ | (10-9 M⊙yr-1) | |||||
| Anon 1 | LAMOST | 2014-11-05 | -3.90.2 | 3.9 | — | — | WTTS |
| BS Tau | HCT | 2021-01-17 | -31.11.2 | 33.12.0 | 0.0110.001 | 2.17 | CTTS |
| 2021-09-06 | -35.11.3 | ||||||
| CIDA 1 | LAMOST | 2012-01-04 | -141.934.3 | 205.5 | 0.004 | 0.70 | CTTS |
| 2011-12-18 | -224.821.6 | ||||||
| 2012-01-13 | -133.316.4 | ||||||
| 2012-01-22 | -97.77.3 | ||||||
| 2013-02-08 | -271.718.9 | ||||||
| 2014-01-25 | -242.128.2 | ||||||
| 2015-10-13 | -231.526.7 | ||||||
| 2015-12-28 | -273.247.6 | ||||||
| 2016-12-31 | -232.911.9 | ||||||
| CW Tau | LAMOST | 2014-01-25 | -87.12.3 | 126.8 | 0.54 | 107.69 | CTTS |
| 2014-01-30 | -232.27.2 | ||||||
| 2014-11-05 | -60.41.8 | ||||||
| FM Tau | LAMOST | 2014-01-30 | -112.83.6 | 118.8 | 0.083 | 16.75 | |
| HCT | 2021-01-17 | -100.73.7 | |||||
| 2021-09-06 | -142.94.2 | ||||||
| FO Tau | LAMOST | 2011-12-18 | -136.49.3 | 143.5 | 0.059 | 11.68 | CTTS |
| 2012-01-04 | -172.713.4 | ||||||
| 2012-01-13 | -135.99.4 | ||||||
| 2013-02-08 | -135.29.3 | ||||||
| 2014-01-25 | -137.47.5 | ||||||
| 2014-01-30 | -158.211.2 | ||||||
| HD283782 | HCT | 2021-09-06 | -1.90.1 | 1.9 | — | — | WTTS |
| V773 Tau | LAMOST | 2014-11-05 | -3.20.1 | 2.5 | — | — | WTTS |
| HCT | 2021-09-06 | -1.90.1 | |||||
| V836 Tau | HCT | 2021-01-17 | -8.90.5 | 13.3 | 0.013 | 2.60 | CTTS |
| 2021-09-06 | -17.60.5 | ||||||
| J041411.88 | LAMOST | 2012-01-22 | -319.435.6 | 357.4 | 0.002 | 0.49 | CTTS |
| 2012-01-23 | -404.6138.3 | ||||||
| 2015-12-28 | -348.269.7 |
Note. — Column 5 represents the average value of from multi-epoch observations as listed in column 4 and the uncertainties denote the range of from multi-epoch observations. The uncertainties in Lacc and listed in columns 6 and 7, respectively, represent the range in these values due to the range in values.